You need to test H0: μ=10 against H1: μ>10. The test statistic was found to be,...

  1. You need to test H0: μ=10 against H1: μ>10. The test statistic was found to be, Ztest=1.72. The P value of the test should be:                                                                                                                                                                

  1. 0.0427
  2. 0.9573
  3. 0.0854

  1. You need to test H0: μ=100 against H1: μ<100. The P value of the test was found to be 0.0001. A possible 95% confidence interval is:                                                                                                                                                    

  1. -∞, 99.5
  2. -∞, 78.6
  3. 110.3, ∞

  1. A 95% upper confidence interval for the tensile strength of 0.05 millimeter (mm) Sisal fiber in Megapascals (MPa) was found to be 601.3, ∞. The correct interpretation of this confidence interval is:   

  1. There is a 95% probability for the true tensile strength of 0.05 mm Sisal fiber exceeds 600 MP
  2. 95% of 0.05 mm Sisal fiber has tensile strengths greater than 601.3 MPa
  3. 95% of 0.05 mm Sisal fiber has tensile strengths lesser than 601.3 MPa

  1. You need to test H0: β=0     vs     H1: β>0 for simple linear regression using a sample of 15 data points, where β is the slope. The test statistic was found to be ttest=2.03. The P value of the test should be:            

  1. 0.0317
  2. 0.0309
  3. 0.9691
Answer #1

1) Answer: a) 0.0427 Explanation: P-value = P(Z > 1.72) = P(O<Z<) - P(0<Z<1.72) = 0.5-0.4573 P-value = 0.0427 (From normal ta